OpenStudy (anonymous):

The volume of the room is 40 1/4 cube meter, and the height is 2 1/2 meter. Find the area of the base.

OpenStudy (aravindg):

volume= base area *height

OpenStudy (aravindg):

can u do it now ?

OpenStudy (anonymous):

so 40 1/4 / 2 1/2 = 161/4/ / 5/2 = 161/2 * 1/5= 16 1/10 Right? :))

OpenStudy (aravindg):

yep !! well done ! \[\checkmark \]

OpenStudy (anonymous):

thank you very much :)